Let λ ∈ R be an eigenvalue of an orthogonal matrix A. Show that λ = ±1.
(Hint: consider the norm of Av, where v is an eigenvector of A associated with the
eigenvalue λ.)

Also, find diagonal orthogonal matrices B, C such that 1 is an eigenvalue of B
and −1 is an eigenvalue of C.
1
Expert's answer
2020-03-24T10:41:54-0400

A square matrix "A" is said to be orthogonal if "AA'=I" "=A'A"

where "A'=" Transpose of "A" and "I=" Identity matrix.

Suppose "\lambda \in R" be a eigenvalue of "A" .

Then there exists a non zero eigenvector "X" such that

"AX=\lambda X" "......(1)"

Taking transpose of both sides of the above equality, we get

"(AX)'=(\lambda X)'"

"\implies" "X'A'=X' \lambda"

"\implies" "X'A'=\lambda X'"

Multiplying both sides by "AX" we get,

"X'A'AX=\lambda X'AX"

"\implies""X'X=\lambda X' \lambda X" [from equation (1)]

"\implies X'X={ \lambda }^2 X'X"

"\implies (1- { \lambda}^2)X'X=0"

Since ,"X \neq0 \implies X'X\neq0."

Therefore,"(1-\lambda^2)=0"

"\implies \lambda^2=1"

Hence "\lambda =1,-1" .

"(Proved)" .

Let "B=" "\begin{pmatrix}
 1&0\\
 0&1
\end{pmatrix}" which is a diagonal matrix and "BB'=B'B=I"

Hence ,"B" is an othogonal matrix ,whose eigen values are "+1, +1."

Let "C=\begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix}" which is a diagonal matrix and "CC'=C'C=I"

Hence ,"C" is an orthogonal matrix, whose eigen values are "-1,-1." .